OpenStudy (anonymous):

HELP! One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

humans = 2x +1 x horses = 4y + 1 y 74 = x + y 196 = 2x + 4y using variables x for humans and y for horses do you know how to solve these two equations algebrically?

OpenStudy (anonymous):

thnk you sista, and naw not really girl

OpenStudy (anonymous):

NP... do you understand how I created the equations?

OpenStudy (anonymous):

yes i think lol

OpenStudy (anonymous):

ok lets make sure = 74 heads we know each human and horse have one of each; so 74=x+y

OpenStudy (anonymous):

196 legs = humans have two (2x) and horses have four (4y) 196 = 2x+ 4y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):